Minutes of the Trigger Operation Meeting (May 30, 2006)
Attendees: Arik Kreisel, Selina Li, Rahmat Rahmat, Dan Walker
- PEP and BaBar have been doing really well for the last week, running at around 1x10^34 and breaking some shift records. Trigger has no operational issues.
- People are reminded to have their IR2 training (and receive a white/silver star for their badge) if they want to help out others who work at the detector hall. Note that the IR2 hall training is optional to our Trigger system worker since our work can be done at the EH and teststand and no need to work on or around the detector. However, please read our safety guide.
- Dan looked at the EMT deadtime in some periods in March thru May. Overall, it looks like the EMT contribution to the deadtime is about 0.02 % in May (after Chris O'Grady fixed a bug in the ROM), as compared to 0.05% in March. See some EMT deadtime plots here. Dan will make a comparison to the running condition last October before the shutdown.
- Because of the ZPD damage we saw every once a while, we need to filter out ZPD damaged events in the Digi (reconstruction) filter after the ICHEP cut-off date.
- We will disassemble the old DCT crate some time in June. Details will be discussed in our next TRG operation meeting when more TRG experts are back in town.
- The two physics_l1passthru runs taken on May 24th are run 65090 and 65091. The second run (65091) has a lumi dip and the run paused twice according to Katherine George. Arik will use these runs to study the L3 tracking performance.
Data Quality and Monitoring
This page is created by Selina Li (firstname.lastname@example.org); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.